Reports 1-1 of 1 Clear search Modify search
CRY (Cryostat IX)
shinji.miyoki - 15:43 Saturday 25 November 2023 (27720) Print this report
One od Cryocooler 50K Head showed strange temp enhencement

About two days ago, the temp of REF3_50K_HEAD showed ~ 10K enhancement up to ~92K from 82K, then the vacuum level at IXA became 5x10^-6 to 1x10^-5Pa.

The temp at the top of the outer shield in IXC is around 167k which is comparable with the temp of EXC's same position (165K). Not only at this position but also at other positions, temps are around 170K which is a critical temp for H2O.

Fig.3 shows the Qmass data. N2 is the main increasing component. H2O is constant.

Fig.4 shows the ~4-month trend after starting cryocoolers for 2-layer radiation shields. Both cryocoolers seem to show temp enhancement. However, one 50K Head (REF1) is now on the recovery, while the other(REF3) shows strange oscillation.

Images attached to this report
Comments to this report:
tomotada.akutsu - 8:18 Saturday 25 November 2023 (27721) Print this report

OMG... Can you also check N2 pressure in the compressor for the relevant coolers...???

shinji.miyoki - 15:43 Saturday 25 November 2023 (27722) Print this report

Temps comparison of 2-layer radiation shield of IXC, IYC and EYC. 

For I{XY}C, Almost the same temps for the inner radiation shield, while IYC outer radiation shields have lower temps by ~ 20K.

For {IE}XC, a little bit lower temps in IXC for the inner radiation shield, while almost the same temps for the outer radiation shield.

Judging from these data, the present situation of shield temperatures in IXC seems to be between the situations of EXC and IYC.

Some temps data in IY seem to be strange, for example, 1) IY 80K Bottom REF1,2 should not be 55K but ~ 140K, 2) 80K shield Side (a) should not be 0.

Images attached to this comment
shinji.miyoki - 13:23 Sunday 26 November 2023 (27723) Print this report

[miyoki, yokozawa, uchiyama(remote)]

IXC P53, P55, Water circulater situations. Not so different from those that were taken before this event.

The main difference was the temperature of the He gas tank in Fig2,4, ~80C, while 65C for the corresponding IYC compressors. According to Kimura-san, this temperature could affect the cooling performance. However, ~80C is the same as that before this event.

  • P53(REF1): Fig1-2, P53-3.mov
  • P55(REF3): Fig3-4, P55-3.mov
  • Water circulator: Fig5-7, Water-4.mov

 

Images attached to this comment
Non-image files attached to this comment
shinji.miyoki - 8:22 Tuesday 28 November 2023 (27742) Print this report

The outer shield temps were enhanced by 1K ~ 4K compared with yesterday as Fig.1. The inner shield temps were also enhanced by ~1K.

The 1st head temp of REF3 cryocooler head was ~ 120K So, 30K enhancement within one day. This phenomenon is almost identical to the phenomena in the cryocooler for the BRT side duct shield in EXC in August 2023. In the case of EXC, 30K enhancement within 1 "hour". (As you know, no connection with the 2nd head in the case of cryocooler for the EXC BRT side duct shield, and no one more additional cryocooler(REF1) as REF3 in IXC.), Fortunately, the cryocooler cooling performance was recovered by itself within 13 days as Fig.2 without any of our actions. Fig.2 shows the temp saturation around 120K. The REF3 temps also seemed to saturate around 120K in this morning as Fig.3. We hope the same phenomena (self-recovery) will happen in REF3. 

The vacuum level at IXA is ~ 1.6 x 10^-5 Pa which is 3 times larger than before. N2/H2O partial pressure (10^-5 Pa, and 2x10^-6 Pa, individually) was also enhanced by ~ 3 times. However, 2x10^-6 Pa H2O partial pressure is 2 times larger than that in the healthy EXC.

We are now considering the TMP operation around IXC, assuming a lower risk of TMPs' stop because they are set in much better condition in the center area than in the arms. TMPs at X-2 and GVbsixc can be isolated by GVs remotely for the emergency case. While 2 TMPs at GVbsixc - ICX - IXA cannot be isolated by GVs remotely. 

Images attached to this comment
shinji.miyoki - 8:50 Tuesday 28 November 2023 (27743) Print this report

After performing calibration? by Kimura-san for Q-mass for IXC, IYC, EXC, EYC, and BS around 8:45, H2O partial pressures are,

  • 2x10^-6 Pa at IXC
  • ~1.5x10^-6 Pa at others.

So, no so large differences.

Around 5:30 before an unlock the PRFPMI+RFlock, the IRX:130 ~ 133, IRY 128~130.

Note: Qmasses for cryostats are set near the cryostat bodies, while the BS Qmass is at near the GVbs-ixc.

takashi.uchiyama - 10:37 Tuesday 28 November 2023 (27747) Print this report
2023/11/28

Miyoki, Uchiyama

We restarted TMP at X-3 which is the BS side of GVbsx at 10:05 on November 28.
Images attached to this comment
shinji.miyoki - 15:54 Tuesday 28 November 2023 (27749) Print this report

I closed GV in front of the TMP, because

  • REF3 1st head temp decreased a little bit (122K -> 118K) and the vacuum level was also on the way of recovery.
  • The partial water pressure levels were almost the same among IXC and IYC: ~1.7x10^-6 Pa, even if the TMP was isolated by GV. This means that TMP seems not to be effective to reduce the water at the Q-mass position.  
  • we would like to minimize the risk of TMP stopping during nighttime.
shinji.miyoki - 20:17 Tuesday 28 November 2023 (27754) Print this report

The temp at the REF3 1st cold head (right column / 3rd row in Fig.1) is now decreasing. We hope the cooling power will be recovered as the cryocooler in EXC as in August 2023.

Images attached to this comment
shinji.miyoki - 10:38 Thursday 30 November 2023 (27774) Print this report

The REF3 1st head temp is now around 90K. It is still on the way of reducing.

The cooling speed is roughly 10K/day, On the other hand, the cooling speed in the case of cooling from 300K, is ~20K/day.

Images attached to this comment
shinji.miyoki - 18:59 Friday 01 December 2023 (27784) Print this report

Anyway, according to the opinions of the Sumitomo Heavy Industry, the situation where the 1st Head temp increased up to 120K and the 2nd Head temp is kept around 25K is an impossible phenomenon, and some troubles should happen somewhere, except for the cryocooler heads and compressors themselves.

Then, we considered other possibilities.

  1. The present length of the He gas flexible tube length (20m) is really good for the long operation? 20m consists of two 10m flexible tubes to obtain the best cool temperature at each cold head according to our experiences.
  2. All flexible tubes for 6 cryocoolers for each cryostat are gathered in one closed pit (40cm x 20cm x ~ 5m roughly with iron plate covering). The temperatures of tubes to the compressors are much hotter than the temp of tubes to the cryocooler. There may be an undesired heat coupling in this pit. 
  3. The airflow in these pits can be different among the IYC case (the cryocooler operation is healthy) and others (at least, one of cryocoolers got troubles). In all clean booths for the cryostats, some FFUs beside KOACHs are working. In IYC, a tunnel is set and it spoils the airflow. This results in the air pressure enhancement at the FFU/pit side, and in the airflow from the clean booth area to the machine room through the pit. On the other hand, in other clean booths, the air from the FFUs can go out easily through the entrance area where you should change your wearing to clean suits. So, even as 2), the only IYC tubes in the pit can be cooled by this airflow.

Kimura-san will investigate these issues.

shinji.miyoki - 17:23 Saturday 02 December 2023 (27807) Print this report

IXC REF3 1st head temp seems to stop reducing around 85K. This is 15K higher than other corresponding cryocoolers at EXC and IYC.

In addition, IXC REF1 1st head temps is also around 80K.

So some reason should exist to explain this higher temperature in IXC.

shinji.miyoki - 6:43 Tuesday 05 December 2023 (27831) Print this report

I checked the He gas tube routings for IXC.

Unlike other cryostats (EXC, IYC, EYC), the He gas tubes are not enclosed in the pit. Because of their long lengths, they are partially set in the pit, and partially on the floor. To avoid trampling, metal mech floors are set over them. In this sense, these He gas tubes have enough ventilation, and actually, the temps of these tubes are moderate around 22C~23C. I tried to cool these tubes with two air circulators. However, there seems to be no effect at all, judging from the 18 hours monitoring.

The temp of the 1st head of REF3 cryocooler is now fluctuating around 85.4K, while 68K for EX and IY.

The staff of the Sumitomo HI asked us again the pressure difference which should be less than 1.65 Pa. Otherwise, the cooling property will be spoiled as EXC and EYC cryocoolers for the duct shields.

shinji.miyoki - 11:00 Saturday 09 December 2023 (27903) Print this report

The 1st head temp of the cryocooler for the BS side duct shield in IXC showed ~15K enhancement and recovery around the early morning today.

The vacuum level at BS seems to show no remarkable enhancement.

Temp enhancement in the BS side duct shield seemed to be within ~1K around 90K~100K.

Images attached to this comment
shinji.miyoki - 7:29 Sunday 10 December 2023 (27915) Print this report

The same warm-up happened again on Sunday mid-night.

The temp change of the 1st cryocooler head was a very smooth enhancement like a natural warmup. Did a temporal stop of something happen?

Images attached to this comment
shinji.miyoki - 9:23 Sunday 10 December 2023 (27916) Print this report

According to Kimura-san, any detected errors result in the stop of the cryocooler and no automatic restart even when the error sources are recovered. The water supply system shows no errors according to the remote webcam image. So, water supply system seems to be healthy. 

In this time in IXC, the warming up time by 15K took 20 minutes. While, it took a shorter time of 15 minutes when the cryocoolers stopped because of the power outage in the EX area on October 28th. So in IXC, the cryocooler itself kept working, but lost its cooling power because of something. 

shinji.miyoki - 16:21 Sunday 10 December 2023 (27918) Print this report

The same enhancement happened around 01:20 UTC on September 27th. The enhancement feature, duration (~ 20 minutes) are almost the same with the recent two events.

Also ~ 5K level temporal enhancement at the 1st head of the cryocooler for the BRT side happened in EXC about 53 days ago from now.

They fortunately recovered.

Images attached to this comment
shinji.miyoki - 10:04 Monday 11 December 2023 (27925) Print this report

Around 9:30 JST, it happened again.

I asked Takahashi-masa san to check the compressor situation for the duct shields, then he found that the He gas tank temp reached around 70C which should have been around 50C before.

As I reported before, the compressors for 2-layer radiation shields in IXC also showed ~ 80C on their He gas tanks in compressors. These cryocoolers also have higher 1st head temperature. So, He gas temp seemed to be related of the cooling power reduction of these cryocoolers.

Images attached to this comment
shinji.miyoki - 13:24 Monday 11 December 2023 (27926) Print this report

I added a fan for the compressor for the cryocooler for the BS side duct shield.

Images attached to this comment
shinji.miyoki - 23:13 Wednesday 13 December 2023 (27962) Print this report

[yoshimura, yamaguchi, miyoki]

Two spot coolers were operated in the machine room of the corner station. The hot air from the spot coolers is now transferred to the 2F of the corner station where a water-cooling type cooler is set.

However, the temp reduction seems to be from 28.6 to 27.8C. Not so effective. The humidity is about 36%.

Actually, several dryers that generate hot air are also operated in the tent for the DGS stock in this machine room, in addition to 8 He gas tanks that have ~ 65C temperature.

shinji.miyoki - 7:11 Friday 15 December 2023 (27976) Print this report

In yesterday's setting of spot coolers, they had no air duct extensions to introduce their cooled air near the compressor's position.

On the 14th, Yoshimura-san's team set such extensions in the morning. Owing to these direct cool air injections around the compressors, the 1st Head temperature of the cryocooler for the BS side duct shield seemed to show a step-like temperature reduction of 0.2C. as Fig.1.

Images attached to this comment
shinji.miyoki - 7:58 Monday 18 December 2023 (27993) Print this report

After the cooling of He gas tank in the compressor and the machine room temp reduction, the strange warm-up in IX_50K_REFBS_HEAD seemed to stop as Fig.1

Images attached to this comment
shinji.miyoki - 8:48 Wednesday 20 December 2023 (28018) Print this report

According to the manual of the compressor checked by Kimura-san, it has an interlock system for the temperature of oil for cooling He gas in the He gas tank. If it senses over 93C, the compressor stops cooling/(operation?) for 20 minutes. If the temp can be below this temp within 20 minutes, it can restart.

The ever all of 4 events in the cryocooler for the BS side duct shield showed rapid temp enhancement for ~ 20 minutes and recooling. So this interlock can explain these events.

This story may also explain the temp enhancement in the cryocooler for the BRT side duct shield on 12th August. In this case, the temp of the 1st head of the cryocooler rapidly increased from 80K to 110K for 1 hour and kept 100~110K  for ~ 3 days, then recooling started. However, we don't know why the temp of oil cannot be cooled for several days.

On the other hand, the main warm-up reason of the cryocooler for the EY arm side duct shield was the He gas leak of the compressor after actions.

However, we have no remarkable reason for the warmup of the cryocooler for the 2-layer radiation shields in IXC. At present, we are just suspecting some heat-up conditions around compressors.

Consequently, 3 of 4 strange performances of cryocoolers can be explained, but no clear answer for one. 

nobuhiro.kimura - 22:44 Wednesday 20 December 2023 (28026) Print this report

[Kimura and  Araki (KEK)]
 Temperatures around a thermal switch mounted on the helium gas outlet piping of the compressor of a BS-side duct shield rcryo-cooler (Xfs) was measured.
 This cryo-cooler has a history of repeated shutdowns and recoveries. (K-log 27955)
The display of the contact thermometer used for the measurement was approximately 94°C.
This value exceeded the thermal sensor's setpoint of 93°C. Even taking into account the sensor's own sensitivity error, the compressor was considered to be operating in a critical condition.
 From the description in the instruction manual, the temperature around this thermal switch is estimated to be around 69°C, which is the normal value.
As a temporary cooling device, a circulator was placed to cool the piping around the thermal switch.
We suggest replacing the compressor.

Images attached to this comment
shinji.miyoki - 7:55 Thursday 21 December 2023 (28038) Print this report

The temp of IX_50K_REFBS_HEAD after the latest cooling start. As you know, the four ~15K enhancements for 20 minutes and recooling happened around 27 Sept and 9th  Dec.

  • Remarkable temp enhancement existed before and after the event on 27th Sept by 0.2C.
  • Around the 1st Nov, the temp increased gradually.
  • 0.2C enhancement again before and after the 3 events around 9th Dec.

 

Images attached to this comment
shinji.miyoki - 14:57 Thursday 21 December 2023 (28045) Print this report

[Yoshimura, Yamaguchi, Miyoki]

We put 4 circulator fans around the compressor. In addition, we set the spot cooler to directly inject the 17C air to the back of the compressor for the BS side duct shield.

We hope, more cooling effect is obtained.

Images attached to this comment
nobuhiro.kimura - 8:20 Friday 22 December 2023 (28051) Print this report

[Kimura and Araki (KEK)]
 The temperature of the piping in the compressor for the  duct shield cryo-cooler at BS side (Xfs) was measured.
The results are shown below.
  Row 2 is without circulator.
  Row 3 is one circulator
  Row 4 is one circulator and spot cooler added.

 We confirmed the temperature of the capillary tube and the base of the compressor is reduced by about 6°C due to the enhanced cooling.
As a result, a margin was secured against the thermal switch attached to the gas outlet piping, which is approximately 93°C..

 

20/Dec.

 9:58~10:10 21/Dec. 15:14~15:30 21/Dec.
Items without circulator one circulator Four circulators and one spot cooler added
Helium gas supply (°C) 21.9 21.3 22.7
Helium gas return (°C) 24.3 23.9 24.0
Adsober OUT (°C) 22.5 21.5 22.8
Adsober IN (°C) 21.1 21.0 22.1
Oil separator IN (°C) 20.1 19.6 19.8
Oil separator OUT (°C) 20.9 20.5 21.1
Heat exchanger He IN (IN-1) (°C) 89.1 78.3 83.9
Heat exchanger OIL IN ( IN-2) (°C) 89.2 85.5 89.0
Surface temp. of the main unit (°C) 65 65 55
Capillary tube temp. for oil return (°C) 68.8 73.3 60.7
Gas outlet piping temperature (°C) 93.4 93.5 87.6
Water inlet temp. (°C)  19.7 18.7 18.3
Water outlet temp. (°C) 25,1 24.5 24.3
Psupply (MPa) 2.28~2.32 2.28~2.32 2.30~2.33
Preturn (MPa) 0.648~0.689 0.662~0.738 0.655~0.710
Pdeff (MPa) 1.63~1.63 1.618~1.582 1.645~ 1.620
shinji.miyoki - 7:43 Wednesday 27 December 2023 (28098) Print this report

Finally, the REF1 also showed the temp enhancement from ~80K to ~100K from -8 days to -4 days from today, and it is now decreasing again from -4 days to now. N2 gas component also increased several days ago and decreased now.

In the case of REF3, the enhancement reached up to ~120K, then started decreasing and fluctuating around 85K.

The same phenomena seem to happen in REF1/3 cryocoolers. 

In addition, REF3 1st head temp also start increasing over 100K. We hope it will turn back at around 120K as before.

Images attached to this comment
shinji.miyoki - 15:25 Wednesday 27 December 2023 (28111) Print this report

Although the temps on the surface of the He gas tank in compressors for 2-layer radiation shields were moderate temps around 65K or so, we added one more circulator for each compressor at the side of the He has tank.

However, no remarkable effect on the temperature at the REF3/REF1 cryocooler 1st heads. 106K at REF3 1st head, 92K at REF1 1st head.

Images attached to this comment
shinji.miyoki - 16:31 Wednesday 27 December 2023 (28113) Print this report

I asked Yoshimura-san to change the circulator position for P-55(REF3?) from the side to the top to cool the heat exchanger that has many fins around 16:00. 

The additional fan for P-53(REF1) was set at the heat exchanger side for side flow to it.

shinji.miyoki - 7:42 Thursday 28 December 2023 (28118) Print this report

Just after cooling the heat exchanger for P-55, the  REF3 1st head temp seemed to stay around 107.8K for 1 hour. However, it started increasing again nad reached around 112K.

At present, the temp seemed to decrease and be around 107K.

While the REF1 1st head temp is now around 92K.

Images attached to this comment
Search Help
×

Warning

×